OpenStudy (marcelie):

Find the points on the curve y= 2x^3+3x^2-12x+1 where the tangent is horizontal.

OpenStudy (marcelie):

@Photon336

OpenStudy (photon336):

let's first get the derivative first

OpenStudy (photon336):

\[\frac{ d }{ dx } = (6x^{2}+6x-12)\]

OpenStudy (marcelie):

hmm im thinking it would look like this y' = 12x+6

OpenStudy (photon336):

so you took the second derivative?

OpenStudy (marcelie):

oh shoot i messed up lol okay so we would equal it to 0 for first derivative right ?

OpenStudy (photon336):

let me see

OpenStudy (marcelie):

6x ^2+6x-12= 0 6( x^2+6x-12)= 0 (x-2 ) (x+1 ) =0 x= -1 and x=2

OpenStudy (photon336):

yeah that's what I got. now let's find the equation for that.

OpenStudy (photon336):

actually I got m = -2 and 1

OpenStudy (sshayer):

when x=1 find y again when x=-2 find y

OpenStudy (marcelie):

oh riight sorry lool hmm so we would plug these points back to the original equation right ?

OpenStudy (photon336):

Yeah okay @marcelie I think I got it

OpenStudy (photon336):

Remember when we took the derivative and set it equal to zero?

OpenStudy (photon336):

\[2(-2)^{3}+3(-2)^{2}-12(-2)+1 = 21\] \[2(1)^{3}+3(1)^{2}-12(1)+1 = -6\]

OpenStudy (photon336):

these are the two points where the derivative = 0 \[m(x-x_{0}) = (y-y_{0})\] so by definition m = 0 when x = -2 or x = 1 so this means \[0*(x-x_{0}) = y-21\] and 0*(x-x_{0}) = y+6 re arranging we get y = 21 and y = -6

OpenStudy (photon336):

@marcelie using the site desmos online graphing calculator graph the original function and y = -6 and y = 21

OpenStudy (marcelie):

oh okay so our answers would be y= -6 and 21 ?

OpenStudy (photon336):

yes

OpenStudy (marcelie):

oh okay x)

OpenStudy (sshayer):

points are (1,-6) and (-2,21)